OpenStudy (anonymous):

need help with a related rates problem. ship A is 100km west of ship B at 12:00. Ship A then sails south at 35km/h and ship B sails north at 25km/h. How fast is the distance between the ships changing at 4:00 PM?

OpenStudy (anonymous):

(100)^2 +(35t+45t)^2

OpenStudy (anonymous):

enjoy. i think equation is something like \[(A+B)^2 + 100^2=D^2\] and then \[2(A+B)(A'+B')=2DD'\] or \[(A+B)(A'+B')=DD'\] plug in numbers solve for D'

OpenStudy (anonymous):

\[A=4\times 35=140\] \[B=4\times 25=100\] \[A'=35\] \[B'=25\] \[D=\sqrt{240^2+100^2}\] and that is that